Even if both the partial derivatives of f : R^2 → R
exist at a point (a,b) , the function f may still not be continuous at (a,b).
1
Expert's answer
2020-03-05T13:58:22-0500

Let we have function "f: E\longrightarrow R, E\sub R^n, E \text{ is an open set}."

Theorem. If "f" is differentiable at a point "x_0" f is continuous at a point "x_0".

In case "n\geq 2" if partial derivatives of f exist f may be not differentiable. For example, consider the function "f(x,y)=1 \text{ if } xy=0 \text{ and } f(x,y)=0 \text{ otherwise}." Then we have "\frac{\partial f}{\partial x}(0,0)=\frac{\partial f}{\partial y}(0,0)=0" but function f is not continuous at (0,0):"lim_{(x,y)\longrightarrow (0,0)}=0\neq f(0,0)=1." So function f is not differentiable at (0,0).